# Motivic classes of stacks in finite characteristic and applications to stacks of Higgs bundles

### v1

@@ Abstract.
We define a ring of motivic classes of stacks suitable for symmetric powers in finite characteristic. Let $X$ be a smooth projective curve over a field of arbitrary characteristic. We calculate the motivic classes of the moduli stacks of semistable Higgs bundles on $X$. This recovers results of Fedorov, A. Soibelman and Y. Soibelman in characteristic zero, as well as those of Mozgovoy and Schiffmann for finite fields. We also obtain a simpler formula for the motivic classes of Higgs bundles in the universal $\lambda$-ring quotient using Mellit’s results.
@@ 1. Introduction and main results
In [FSS18] similar formulas were given for motivic volumes of the stacks of Higgs bundles. It would be nice to reprove the formulas of [Sch14] and [MS20] by applying counting measures to motivic formulas. Unfortunately, it is assumed in [FSS18] that characteristic is zero. In this paper, we fix this problem by giving the “corrected” definitions of rings of motivic classes of stacks so that the formulas of [FSS18] become valid in finite characteristic as well. Thus, our main theorem generalizes both [FSS18, Theorem 1.3.3] and [MS20, Theorem 1.1(ii)]. We also obtain simpler formulas for the motivic classes of Higgs bundles in the universal $\lambda$-ring quotient using Mellit’s results, see [Mel18, Mel20a].
@@ 1.1. Motivic classes of stacks
First we define the motivic classes of stacks. Compared to the characteristic zero case, the difference is that we need to add a “surjective and universally injective relation” for the definitions of motivic classes.
@@
Definition 1.2.
@@
    $\mathcal{X}]=[\mathcal{Y}]+[\mathcal{X}-\mathcal{Y}]$ where $\mathcal{Y}$ is a closed substack of $\mathcal{X}$;
@@
    $\mathcal{X}]=[\mathcal{Y}]$ if there is a surjective and universally injective morphism $\mathcal{X}\to\mathcal{Y}$ of stacks over $k$;
@@
    $\mathcal{X}]=[\mathcal{Y}\times\mathbb{A}^{r}_{k}]$ where $\mathcal{X}\to\mathcal{Y}$ is a vector bundle of rank $r$.
The class $\mathcal{X}]$ in $\operatorname{Mot}(k)$ is called the motivic class of the stack $\mathcal{X}$. The class of $\mathbb{A}^{1}_{k}$ in $\operatorname{Mot}(k)$ is denoted by $\mathbb{L}$. We can give a ring structure for $\operatorname{Mot}(k)$ by $\mathcal{X}]\cdot[\mathcal{Y}]\coloneqq[\mathcal{X}\times\mathcal{Y}]$.
@@ 1.3. Explicit formulas for motivic classes
We have a similar result compared to [FSS18, Theorem 1.3.3] for the motivic classes of moduli stacks of semistable Higgs bundles. In particular, for $r\geq 0$ we define classes $H_{r,d}\in\overline{\operatorname{Mot}}(k)$ by complicated but explicit formulas, see Sections 5.4 and 5.5. The following theorem gives an explicit formula for the motivic classes of stacks of Higgs bundles.
@@ 1.4. Mellit’s simplification
We define the motivic zeta function for every reduced quasi-projective scheme $Y$:
$\zeta_{Y}(z)=\sum_{n\geq 0}[Sym^{n}Y]\cdot z^{n}\in 1+z\cdot\operatorname{Mot}(k)[[z]],$
Recall that $X$ is a smooth projective geometrically connected curve over $k$. We now give the definition of the motivic $L$-function, which is a polynomial of degree $2g$, where $g\coloneqq g(X)$ is the genus of $X$, see [FSS18, Proposition 1.3.1]:
@@
Here the sum is over all partitions $\lambda=(\lambda_{1}\geq\ldots\geq\lambda_{i}\geq\ldots)$ and the product is over all boxes of the partition, $a(\square)$ and $l(\square)$ are the arm and the leg of the box respectively, $\langle\lambda,\lambda\rangle\coloneqq\sum_{i}(\lambda^{\prime}_{i})^{2}$ where $\lambda^{\prime}=(\lambda^{\prime}_{1}\geq\ldots\geq\lambda^{\prime}_{i}\geq\ldots)$ is the conjugate partition and $|\lambda|=\sum_{i}\lambda_{i}$.
@@ 1.5. Organization of the article
The paper is organized as follows: in Section 2, we discuss motivic classes of stacks in finite characteristic compared to the characteristic 0 case in [FSS18]; in Section 3, we define the counting measure for the ring of motivic classes; in Section 4, we define the power structures on the rings of motivic classes of schemes and stacks; in Section 5, we calculate the motivic classes of the stacks of semistable Higgs bundles; finally in Section 6, we use Mellit’s results [Mel18, Mel20a] to obtain a simpler formula for the motivic classes of Higgs bundles in the universal $\lambda$-ring quotient and to calculate their volumes in the case when the field is finite.
@@ 2.1. Definitions of the rings of motivic classes of schemes and stacks
In this section, we give basic definitions and results of motivic classes of schemes and stacks. In finite characteristic, we give a new definition for the ring of motivic class of stacks. We recall that the motivic classes of varieties have been studied in [BM15, Göt01, Kap00, LL03, LL04, Mus11]; the motivic classes of stacks have been studied in [BD07, BM15, Eke09, FSS18, Joy07, Toë05]; the motivic functions have been studied in [FSS18, KS08].
@@
    $\mathcal{X}]=[\mathcal{Y}]+[\mathcal{X}-\mathcal{Y}]$ where $\mathcal{Y}$ is a closed substack of $\mathcal{X}$;
@@
    $\mathcal{X}]=[\mathcal{Y}]$ if there is a surjective and universally injective morphism $\mathcal{X}\to\mathcal{Y}$ of stacks over $k$;
@@
    $\mathcal{X}]=[\mathcal{Y}\times\mathbb{A}^{r}_{k}]$ where $\mathcal{X}\to\mathcal{Y}$ is a vector bundle of rank $r$.
The class $\mathcal{X}]$ in $\operatorname{Mot}(k)$ is called the motivic class of the stack $\mathcal{X}$. The class of $\mathbb{A}^{1}_{k}$ in $\operatorname{Mot}(k)$ is denoted by $\mathbb{L}$. We can give a ring structure for $\operatorname{Mot}(k)$ by $\mathcal{X}]\cdot[\mathcal{Y}]\coloneqq[\mathcal{X}\times\mathcal{Y}]$.
@@
    We need the affine stabilizer condition in Convention 1.1 since by [Kre99, Lemma 3.5.1, Propositions 3.5.6 and 3.5.9], every such stack has a stratification by global quotients of the form $X/\operatorname{GL}(n)$ with $X$ a scheme. This will be used in the proof of Theorem 2.6;
@@
    Note that $\emptyset]=0$. Note also that for any stack $\mathcal{X}$ we have $\mathcal{X}_{red}]=[\mathcal{X}]$. Indeed, $\mathcal{X}-\mathcal{X}_{red}=\emptyset$ and we can use the first relation of Definition 2.1;
@@
    We will explain why we need to add the second relation in Lemma 2.12;
@@
    With the second relation of Definition 2.1, we can replace the first relation with apparently a weaker one $\mathcal{X}]=\sum_{i=1}^{n}[\mathcal{X}_{i}]$ if $\mathcal{X}$ is the disjoint union of substacks $\mathcal{X}_{i}$;
@@
    The second relation is non-trivial by the following example. Fix a perfect field $k$ where $\text{char}(k)=p\neq 0,2$. For an elliptic curve $E$ over $k$, denote by $E^{(p)}$ the Frobenius twist of $E$. The relative Frobenius $E\to E^{(p)}$ is surjective and by [SP18, Tag0CCB], it is universally injective, so we have
@@
    Recall the $j$-invariant of $E$ from [Har77, Chapter 4, Section 4]. For any elliptic curve $E$ given in Legendre form $y^{2}=x(x-1)(x-\lambda)$, $E^{(p)}$ is of Legendre form $y^{2}=x(x-1)(x-\lambda^{p})$, then
    $j(E^{(p)})=2^{8}\frac{\left(\left(\lambda^{p}\right)^{2}-\lambda^{p}+1\right)^{3}}{\left(\lambda^{p}\right)^{2}(\lambda^{p}-1)^{2}}=\left(2^{8}\right)^{p}\frac{\left(\lambda^{2p}-\lambda^{p}+1\right)^{3}}{\lambda^{2p}(\lambda^{p}-1)^{2}}=(j(E))^{p}.$
    By [Har77, Chapter 4, Theorem 4.1(b,c)], in general $E^{(p)}\not\simeq E$ unless $k=\mathbb{Z}/p\mathbb{Z}$;
@@
    [X]=[Y]+[XY][X]=[Y]+[X-Y] where $Y$ is a closed subscheme of $X$;
@@
The class $[X]$ in $\operatorname{Mot}_{sch}(k)$ is called the motivic class of the scheme $X$. The class of $\mathbb{A}^{1}_{k}$ in $\operatorname{Mot}_{sch}(k)$ is also denoted by $\mathbb{L}$ by abuse of notation. Similarly, we can give a ring structure for $\operatorname{Mot}_{sch}(k)$ by $X]\cdot[Y]\coloneqq[X\times Y]$.
@@
    In [Mus11, Section 7.2], Mustaţă defines the ring of motivic classes of varieties (i.e., reduced schemes of finite type) $\widetilde{K}_{0}(\text{Var}/k)$ similarly to Definition 2.3 with the only difference that the multiplication is given by $X]\cdot[Y]\coloneqq[(X\times Y)_{red}]$ (the reason for this modification is that the product of two reduced schemes is not always reduced). This ring is, however, isomorphic to $\operatorname{Mot}_{sch}(k)$ because $X]=[X_{red}]$ for a scheme $X$;
@@
    Mustaţă also defines the ring of motivic classes of quasi-projective varieties $\widetilde{K}^{qpr}_{0}(\text{Var}/k)$. This ring is also isomorphic to the previous one, see Lemma 2.17;
@@
    $\operatorname{GL}(n)]=(\mathbb{L}^{n}-1)(\mathbb{L}^{n}-\mathbb{L})\cdots(\mathbb{L}^{n}-\mathbb{L}^{n-1})$;
@@
    For a stack $\mathcal{X}$, we have $\mathcal{X}/\operatorname{GL}(n)]=[\mathcal{X}]/[\operatorname{GL}(n)]$;
@@
$B((X-Y)\times\operatorname{BGL}(n))=([X]-[Y])/[\operatorname{GL}(n)],$

### v2

@@ Abstract.
We define a ring of motivic classes of stacks suitable for symmetric powers in finite characteristic. Let $X$ be a smooth projective curve over a field of arbitrary characteristic. We calculate the motivic classes of the moduli stacks of semistable Higgs bundles on $X$. This recovers results of Fedorov, A. Soibelman and Y. Soibelman in characteristic zero, as well as those of Mozgovoy and Schiffmann for finite fields. We also obtain a simpler formula for the motivic classes of the stacks of Higgs bundles in the universal $\lambda$-ring quotient using Mellit’s results.
@@ 1. Introduction and main results
In [FSS18] similar formulas were given for motivic volumes of the stacks of Higgs bundles. It would be nice to reprove the formulas of [Sch14] and [MS20] by applying counting measures to motivic formulas. Unfortunately, it is assumed in [FSS18] that characteristic is zero. In this paper, we fix this problem by giving the “corrected” definitions of rings of motivic classes of stacks so that the formulas of [FSS18] become valid in finite characteristic as well. Thus, our main theorem generalizes both [FSS18, Theorem 1.3.3] and [MS20, Theorem 1.1(ii)]. We also obtain simpler formulas for the motivic classes of the stacks of Higgs bundles in the universal $\lambda$-ring quotient using Mellit’s results, see [Mel18, Mel20a].
@@ 1.1. Motivic classes of stacks
Compared to the characteristic zero case, the difference is that we need to add a “surjective and universally injective relation” for the definitions of motivic classes.
@@
Recall from [SP18, Tag04XG] that a point of a stack $\mathcal{X}$ is an equivalence class of morphisms $\operatorname{Spec}k\to\mathcal{X}$, where $k$ is a field. We denote by $|\mathcal{X}|$ the set of points of $\mathcal{X}$. A morphism $f:\mathcal{X}\to\mathcal{Y}$ of $k$-stacks is called universally injective if for every morphism $\mathcal{Y}^{\prime}\to\mathcal{Y}$, the induced map $|\mathcal{Y}^{\prime}\times_{\mathcal{Y}}\mathcal{X}|\to|\mathcal{Y}^{\prime}|$ is injective.
Definition 1.2 (Definition 2.1).
@@
    $\mathcal{X}]=[\mathcal{Y}]+[\mathcal{X}-\mathcal{Y}]$ where $\mathcal{Y}$ is a closed substack of $\mathcal{X}$.
@@
    $\mathcal{X}]=[\mathcal{Y}]$ if there is a surjective and universally injective morphism $\mathcal{X}\to\mathcal{Y}$ of stacks over $k$.
@@
    $\mathcal{X}]=[\mathcal{Y}\times_{k}\mathbb{A}^{r}_{k}]$ where $\mathcal{X}\to\mathcal{Y}$ is a vector bundle of rank $r$.
The class $\mathcal{X}]$ in $\operatorname{Mot}(k)$ is called the motivic class of the stack $\mathcal{X}$. The class of $\mathbb{A}^{1}_{k}$ in $\operatorname{Mot}(k)$ is denoted by $\mathbb{L}$. We can give a ring structure for $\operatorname{Mot}(k)$ by $\mathcal{X}]\cdot[\mathcal{Y}]\coloneqq[\mathcal{X}\times_{k}\mathcal{Y}]$.
@@
1.2. Motivic zeta-function and plethystic exponents
We define the motivic zeta-function for every reduced quasi-projective scheme $Y$:
$\zeta_{Y}(z)\coloneqq\sum_{n\geq 0}[Sym^{n}Y]\cdot z^{n}\in 1+z\cdot\operatorname{Mot}(k)[[z]],$
We will see in Section 2.2 that $\overline{\operatorname{Mot}}(k)$ is topologically generated by the classes $\mathbb{L}^{-n}[Y]$ where $Y$ is a reduced quasi-projective scheme. We will define $\operatorname{Exp}$ to be the unique continuous homomorphism
$\operatorname{Exp}\colon\overline{\operatorname{Mot}}[[z]]^{+}\to(1+\overline{\operatorname{Mot}}[[z]]^{+})^{\times}$
such that
$\operatorname{Exp}(\mathbb{L}^{-n}[Y]z)=\zeta_{Y}(\mathbb{L}^{-n}z).$
$\operatorname{Exp}$ is an isomorphism, so we denote by $\operatorname{Log}\coloneqq\operatorname{Exp}^{-1}$ its inverse isomorphism.
The following theorem is a generalization of [FSS18, Lemma 3.8.2] to arbitrary characteristic.
Theorem 1.3 (Theorem 4.6).
Assume that $M$ is a scheme, $\mathcal{A}_{0}=\operatorname{Spec}k$ and $\mathcal{A}_{1},\mathcal{A}_{2},\ldots,\mathcal{A}_{n},\ldots$ are stacks. Put $A(z)=[\mathcal{A}_{0}]+[\mathcal{A}_{1}]z+[\mathcal{A}_{2}]z^{2}+\cdots+[\mathcal{A}_{n}]z^{n}+\ldots$. Then we have
$\operatorname{Exp}(m\operatorname{Log}(A(z)))=1+\sum_{k=1}^{\infty}\left\{\sum_{k_{i}\colon\sum ik_{i}=k}\left[\left(\left(M^{\sum_{i}k_{i}}\backslash\Delta\right)\times\prod_{i}\mathcal{A}_{i}^{k_{i}}\right)/\prod_{i}{S_{k_{i}}}\right]\right\}\cdot z^{k},$
where $\Delta$ is the “large diagonal” in $M^{\sum_{i}k_{i}}$ which consists of $\left(\sum_{i}k_{i}\right)$-tuples of points of $M$ with at least two coinciding ones.
This theorem is crucial for calculating motivic classes of various stacks such as stacks of Higgs bundles. In finite characteristic, this theorem is not valid without the “surjective and universal injective relations” we have added to the definition of the ring of motivic classes.
1.3. Counting measure
Fix a prime power $q$. If $\mathcal{X}$ is a stack of finite type over the finite field $\mathbb{F}_{q}$, we denote by $|\mathcal{X}(\mathbb{F}_{q})|$ the volume of the groupoid $\mathcal{X}(\mathbb{F}_{q})$.
We denote by $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})$ the image of $\operatorname{Mot}(\mathbb{F}_{q})$ in $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$. We want to pass from pre-$\lambda$-rings to $\lambda$-rings (see Section 2.6). We denote by $\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})$ the image of the composition
We have the following ring homomorphism, called the counting measure.
Theorem 1.4 (Theorem 3.2).
  1. (i)
    We have a unique ring homomorphism
    $\colon\operatorname{Mot}(\mathbb{F}_{q})\to\mathbb{Q}$
  2. (ii)
    For any $A\in\operatorname{Mot}(\mathbb{F}_{q})$, the counting measure #(A)\#(A) depends only on the image of $A$ in $\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})$.
@@ 1.5. Explicit formulas for motivic classes
We have a similar result compared to [FSS18, Theorem 1.3.3] for the motivic classes of moduli stacks of semistable Higgs bundles. In particular, for $r\geq 0$ we define classes $H_{r,d}\in\overline{\operatorname{Mot}}(k)$ by complicated but explicit formulas, see Sections 5.4 and 5.5. The following theorem gives an explicit formula for the motivic classes of the stacks of Higgs bundles.
@@ 1.6. Mellit’s simplification
Recall that $X$ is a smooth projective geometrically connected curve over $k$. We now give the definition of the motivic $L$-function, which is a polynomial of degree $2g$, where $g\coloneqq g(X)$ is the genus of $X$, see [Kap00, Theorem 1.1.9] and [FSS18, Proposition 1.3.1]:
@@
Here the sum is over all partitions $\lambda=(\lambda_{1}\geq\ldots\geq\lambda_{i}\geq\ldots)$ and the product is over all boxes of the partition, $a(\square)$ and $l(\square)$ are the arm and the leg of the box respectively, $\langle\lambda,\lambda\rangle\coloneqq\sum_{i}(\lambda^{\prime}_{i})^{2}$ where $\lambda^{\prime}=(\lambda^{\prime}_{1}\geq\ldots\geq\lambda^{\prime}_{i}\geq\ldots)$ is the conjugate partition and $|\lambda|\coloneqq\sum_{i}\lambda_{i}$. Note that we expand the denominators in positive powers of $z$.
@@
Using the theorem above, we can derive the counting results of Schiffmann and Mozgovoy ([Sch14, MS20]), see Theorem 6.3.
@@ 1.7. Organization of the article
The paper is organized as follows: in Section 2, we define and study the motivic classes of stacks in arbitrary characteristic; in Section 3, we define the counting measure for the ring of motivic classes and use the étale realization to show that the counting measure only depends on the image of the motivic class in the completed $\lambda$-ring; in Section 4, we define the power structures on the rings of motivic classes of schemes and stacks; in Section 5, we calculate the motivic classes of the stacks of semistable Higgs bundles; finally in Section 6, we use Mellit’s results [Mel18, Mel20a] to obtain a simpler formula for the motivic classes of stacks of Higgs bundles in the universal $\lambda$-ring quotient and to calculate their volumes in the case when the field is finite.
@@ 2.1. Definitions of the rings of motivic classes of schemes and stacks
In this section, we give basic definitions and results about motivic classes of schemes and stacks. Our definition is new in finite characteristic. We recall that the motivic classes of varieties have been studied in [BM15, Göt01, Kap00, LL03, LL04, Mus11]; the motivic classes of stacks have been studied in [BD07, BM15, Eke09, FSS18, Joy07, Toë05]; the motivic functions have been studied in [FSS18, KS08].
@@
    $\mathcal{X}]=[\mathcal{Y}]+[\mathcal{X}-\mathcal{Y}]$ where $\mathcal{Y}$ is a closed substack of $\mathcal{X}$.
@@
    $\mathcal{X}]=[\mathcal{Y}]$ if there is a surjective and universally injective morphism $\mathcal{X}\to\mathcal{Y}$ of stacks over $k$.
@@
    $\mathcal{X}]=[\mathcal{Y}\times_{k}\mathbb{A}^{r}_{k}]$ where $\mathcal{X}\to\mathcal{Y}$ is a vector bundle of rank $r$.
The class $\mathcal{X}]$ in $\operatorname{Mot}(k)$ is called the motivic class of the stack $\mathcal{X}$. The class of $\mathbb{A}^{1}_{k}$ in $\operatorname{Mot}(k)$ is denoted by $\mathbb{L}$. We can give a ring structure for $\operatorname{Mot}(k)$ by $\mathcal{X}]\cdot[\mathcal{Y}]\coloneqq[\mathcal{X}\times_{k}\mathcal{Y}]$.
@@
    We need the affine stabilizer condition in Convention 1.1 since by [Kre99, Lemma 3.5.1, Propositions 3.5.6 and 3.5.9], every such stack has a stratification by global quotients of the form $X/\operatorname{GL}(n)$ with $X$ a scheme. This will be used in the proof of Theorem 2.6.
@@
    Note that $\emptyset]=0$. Note also that for any stack $\mathcal{X}$ we have $\mathcal{X}_{red}]=[\mathcal{X}]$. Indeed, $\mathcal{X}-\mathcal{X}_{red}=\emptyset$ and we can use the first relation of Definition 2.1.
@@
    We will explain why we need to add the second relation in Lemma 2.12.
@@
    With the second relation of Definition 2.1, we can replace the first relation with apparently a weaker one $\mathcal{X}]=\sum_{i=1}^{n}[\mathcal{X}_{i}]$ if $\mathcal{X}$ is the disjoint union of substacks $\mathcal{X}_{i}$.
@@
    The second relation is non-trivial by the following example. Fix a perfect field $k$ where $p\coloneqq\text{char}(k)\neq 0,2$. For an elliptic curve $E$ over $k$, denote by $E^{(p)}$ the Frobenius twist of $E$. The relative Frobenius $E\to E^{(p)}$ is surjective and by [SP18, Tag0CCB], it is universally injective, so we have
@@
    Recall the $j$-invariant of $E$ from [Har77, Chapter 4, Section 4]. For any elliptic curve $E$ given in Legendre form as $y^{2}=x(x-1)(x-\lambda)$, $E^{(p)}$ is of Legendre form $y^{2}=x(x-1)(x-\lambda^{p})$, thus
    $j(E^{(p)})=2^{8}\frac{\left(\left(\lambda^{p}\right)^{2}-\lambda^{p}+1\right)^{3}}{\left(\lambda^{p}\right)^{2}(\lambda^{p}-1)^{2}}=(j(E))^{p}.$
    By [Har77, Chapter 4, Theorem 4.1(b,c)], in general $E^{(p)}\not\simeq E$ unless $k=\mathbb{Z}/p\mathbb{Z}$.
@@
    [X]=[Y]+[XY][X]=[Y]+[X-Y] where $Y$ is a closed subscheme of $X$.
@@
The class $[X]$ in $\operatorname{Mot}_{sch}(k)$ is called the motivic class of the scheme $X$. The class of $\mathbb{A}^{1}_{k}$ in $\operatorname{Mot}_{sch}(k)$ is also denoted by $\mathbb{L}$ by abuse of notation. Similarly, we can give a ring structure for $\operatorname{Mot}_{sch}(k)$ by $X]\cdot[Y]\coloneqq[X\times_{k}Y]$.
@@
    In [Mus11, Section 7.2], Mustaţă defines the ring of motivic classes of varieties (i.e., reduced schemes of finite type) $\widetilde{K}_{0}(\text{Var}/k)$ similarly to Definition 2.3 with the only difference that the multiplication is given by $X]\cdot[Y]\coloneqq[(X\times_{k}Y)_{red}]$ (the reason for this modification is that the product of two reduced schemes is not always reduced). This ring is, however, isomorphic to $\operatorname{Mot}_{sch}(k)$ because $X]=[X_{red}]$ for a scheme $X$.
@@
    Mustaţă also defines the ring of motivic classes of quasi-projective varieties $\widetilde{K}^{qpr}_{0}(\text{Var}/k)$. This ring is also isomorphic to the previous one, see Lemma 2.17.
@@
    $\operatorname{GL}(n)]=(\mathbb{L}^{n}-1)(\mathbb{L}^{n}-\mathbb{L})\cdots(\mathbb{L}^{n}-\mathbb{L}^{n-1})$.
@@
    For a stack $\mathcal{X}$, we have $\mathcal{X}/\operatorname{GL}(n)]=[\mathcal{X}]/[\operatorname{GL}(n)]$.
@@
$B((X-Y)\times_{k}\operatorname{BGL}(n))=([X]-[Y])/[\operatorname{GL}(n)],$